OpenStudy (emmboo959782397):

simplify square root of 10 - 4 square root of 10

Mehek (mehek14):

\[\sqrt{10}-4\sqrt{10}\] Does it look like that?

OpenStudy (emmboo959782397):

yes

OpenStudy (oraclethinktank):

I believe it's the square root of 10 minus 4 times the square root of 10.

Mehek (mehek14):

It's the same as subtracting 1 - 4 but keeping \[\sqrt{10}\] at the end

OpenStudy (emmboo959782397):

so -3 square root of 10?

Mehek (mehek14):

Yeah \[-3\sqrt{10}\]

OpenStudy (emmboo959782397):

thank you so much

Mehek (mehek14):

You're welcome :)